FT MICROWAVE SPECTROSCOPIC STUDY OF (CH_{3})_{3}N⋯HCN⋯HCN

\underline{M. CRADDOCK}, C. S. BRAUER, and K. R. LEOPOLD, Department of Chemistry, University of Minnesota, Minneapolis, MN 55455; S. KRON,M. DVORAK, Department of Chemistry, St. Cloud State University, St. Cloud, MN 56301.

Rotational spectra of the hydrogen-bonded trimer (CH_{3})_{3}N⋯HCN⋯HCN and10 isotopically substituted derivatives have been observed using Fourier transform microwave spectroscopy. The complex is a symmetric top with the atomic arrangements indicated above. A central focus of this work is to test the effect of microsolvation in promoting proton transfer across a strong hydrogen bond. Analysis of the rotational and quadrupole coupling constants will be presented, along with ab initio binding energy and geometry optimization results. Preliminary structural analysis indicates an N⋯H hydrogen bond contraction of approximately 0.08 Å relative to the dimer (CH_{3})_{3}N⋯HCN. Changes in the secondary hydrogen bond between the two HCN subunits lie within the experimental uncertainty for the N⋯H distance in HCN dimer. The effect of increasing the basicity of the amine in amine-HCN systems will be discussed in the context of previous work on the closely related system H_{3}N⋯HCN⋯HCN^{a}.

aR. S. Ruoff, T. Emilsson, C. Chuang,T. D. Klots,H. S. Gutowsky J. Chem. Phys. \underline{93}(9), 6363, (1990).